I am struggling with graphing and slopes I need to graph the line passing through (4,-4) with a slope of 4/3
Answer by xachu4u(10)

Easy question:

Any line is given generally by : y = mx + c
where m is the slope and c is the intercept.
Given m = 4/3.
=>y = 4/3 x + c
now it passes through, (4,-4)
=>x = 4, y = -4
Substitute and get the intercept c.

Once you get the line , draw the graph.
That I assume you know.
Happy.....